# A semi-parametric marginalized dynamic conditional correlation framework

### v3

@@ 4 Simulation study
The results of the simulation study then provide support to the use of the two-step AL based estimation method which is able to produce accurate parameter estimates and portfolio risk forecasts, for both spherical and non-spherical error distributions.
@@ 5 Empirical Study
The performance of the DCC-AL model in portfolio risk forecasting has been assessed via an empirical study on a multivariate time series of US stock returns. After providing a short description of the data and describing the forecasting design in Section 5.1, we assess the risk forecasting performance of the proposed models for an equally weighted portfolio in Section 5.2. This choice does not imply any loss of generality since our investigation could be easily replicated under alternative portfolio configurations. Our preference for the equally weighted scheme is motivated by the robust performance of this simple allocation rule that, in many instances, has been found to be competitive with more sophisticated benchmarks (DeMiguel et al., 2007). In addition, in Section 5.3 we also evaluate the effectiveness of the proposed framework in generating minimum risk portfolios.
@@ 5.1 Equally weighted portfolio risk forecasting
Several existing univariate semi-parametric models are selected as benchmarks for comparison. First, we consider the recently proposed ES-CAViaR model (Taylor, 2019) with Symmetric Absolute Value (SAV) and Indirect GARCH (IG) specifications for the quantile regression equation. Regarding the ES specification, to facilitate the performance comparison with the DCC-AL, we choose the model with the constant multiplicative ES to VaR factor $(1+\text{exp}(\gamma_{0}))$, which is the factor that we use in the proposed DCC-AL model.
@@
For all the DCC benchmarks, in order to guarantee a fair comparison with the DCC-AL model, the fitted univariate volatility specifications are given by GARCH(1,1) models.
In Figure (2), it is interesting to note that the estimated correlation parameters in the DCC-AL model vary over the forecasting period, reacting to changes in the underlying market volatility level. In particular, the estimated value of $a$ is characterized by a positive trend originating at the outbreak of the pandemic COVID-19 crisis. An opposite behaviour is observed for $b$.
Figure 2: Plots of estimated correlation parameters $a$ and $b$ of the DCC-AL model. The vertical lines denote the outbreak of the COVID-19 emergency (as officially declared by the World Health Organization on 30 January 2020). Whole forecasting sample (1213 days).
Table 2 compares the DCC-AL estimates of correlation coefficients with those obtained for the benchmark DCC specifications considered, reporting the average estimates of correlation parameters $a$ and $b$ across all forecasting steps. Here, consistent with previous findings in the literature, standard DCC estimators tend to oversmooth (estimated $b$ very close to 1) correlations, while this is not the case for DCC-AL models. The DCC-clik estimates stay in between. It is here worth noting that the DCC-clik and DCC-AL models are based on different first stage volatility estimators: QML-GARCH for DCC-clik, and ES-CAViaR-IG as in Equation (6) for DCC-AL. In addition, Figure 3 shows the cross-sectional averages of the estimated conditional correlations (calculated based on $P_{t}$ in Equation (11)) across the forecasting period. As can be seen, the dynamics of the estimated conditional correlations between DCC-AL and DCC-clik are relatively close to each other.
Table 2: Average estimates (across time) of correlation dynamic parameters ($a$ and $b$) from different models and estimation methods. Whole forecasting sample (1213 days).
DCC-AL DCC-QML DCC-t DCC-clik
@@
Figure 3: Cross-sectional averages of the estimated conditional correlation for DCC-QML, DCC-t, DCC-clik and DCC-AL across the whole forecasting sample (1213 days).
The average of estimated first stage univariate modeling coefficients, across all assets and forecasting steps, is reported in Table 3. GARCH-$t$ and QML-GARCH, used in DCC-t and DCC-QML respectively, appear to be less reactive (smaller $\alpha$ estimates) to past shocks compared to the ES-CAViaR-IG, used in the DCC-AL model.
Table 3: Average estimates (across times and assets) of the first stage univariate modeling parameters ($\alpha$ and $\beta$) from different models and estimation methods across the full forecasting sample (1213 days). For ES-CAVIAR-IG, we report $\alpha=\alpha^{(q)}/q^{2}$.
@@ 5.2 Evaluation of VaR and ES forecasts
Assuming equally weighted portfolio returns, one-step-ahead forecasts of daily VaR and ES from the proposed DCC-AL model are produced by using Equation (14). In this section, these forecasts are compared with the ones from the competing models presented in Section 5.1.
@@
$\sum_{t=T_{\text{in}}+1}^{T_{\text{in}}+T_{\text{out}}}(r_{t}-\widehat{Q}_{t})(\alpha-I(r_{t}\leq\widehat{Q}_{t}))\,\,,$ (18)
@@
$S_{t}(r_{t},\widehat{Q}_{t},\widehat{\text{ES}}_{t})=-\text{log}\left(\frac{\alpha-1}{\widehat{\text{ES}}_{t}}\right)-{\frac{(r_{t}-\widehat{Q}_{t})(\alpha-I(r_{t}\leq\widehat{Q}_{t}))}{\alpha\widehat{\text{ES}}_{t}}}.$ (19)
@@
In our analysis, we use the joint loss $S=\sum_{t=T_{\text{in}}+1}^{T_{\text{in}}+T_{\text{out}}}S_{t}$ to formally and jointly assess and compare the VaR and ES forecasts from all models.
@@
Figure 4 visualizes the 2.5% portfolio ES forecasts from the DCC-QML, DCC-clik-QML and DCC-AL models for the Jan 2019–Sep 2021 period, using equally weighted Dow Jones returns. In general, we can see that the ES forecasts produced from the DCC-AL model have a long-run behaviour comparable to the ones from the DCC-QML and DCC-clik-QML. However, in the short term, forecasts from these models can be characterized by substantially different dynamic patterns. This is particularly evident in the period immediately following the outbreak of the COVID-19. Inspecting the ES forecasts at the beginning of 2020, we can see that the three models have distinctive behaviours. Comparing to the DCC-QML model, the DCC-AL model is more reactive to the return shocks. This is because of the larger $a$ and $\alpha$ estimates as discussed in Tables 2 and 3. The ES forecasts from the DCC-clik-QML stay in between the ones from DCC-QML and DCC-AL, which is also consistent with the findings from Table 2.
@@
Figure 4: 2.5% ES forecasts from the DCC-QML, DCC-clik-QML and DCC-AL models, using equally weighted Dow Jones returns.
@@
The quantile and joint loss results from the proposed DCC-AL model and other competing models are shown in Table 4. Overall, the results show that the proposed DCC-AL model, in comparison to the other models under analysis, generates competitive loss results, which lends evidence on the validity of the proposed semi-parametric framework. Furthermore, the DCC-AL model generates smaller loss values than all other parametric or semi-parametric DCC models that have been taken as benchmarks.
@@
The Model Confidence Set (MCS) of Hansen et al. (2011) is employed to statistically compare the quantile loss (Equation (18)) and joint loss (Equation (19)) values yielded by the different models. A MCS is a set of models, constructed such that it contains the best model with a given level of confidence, selected as either 75% in our paper. The SQ method (see Hansen et al., 2011, for details) is employed to calculate the MCS test statistic. The MCS results, using both quantile and joint loss functions, are also included in Table 4, with red highlighting indicates models that are exlcuded from the 75% MCS. As can be seen, the DCC-AL is included in the MCS both quantile and joint loss test. However, all the DCC models, except the proposed DCC-AL, are excluded from the MCS based on the joint loss.
@@
Table 4: The quantile loss (Equation (18)) and joint loss function (Equation (19)) values for all competing models across the full forecasting sample. Models in red are excluded from the 75% MCS according to the SQ statistic.
Model Quantile loss Joint loss
CARE-SAV 90.4 2,405.4
CARE-IG 89.1 2,392.1
ES-CAViaR-SAV 89.4 2,398.1
ES-CAViaR-IG 87.3 2,369.5
DCC-QML 98.7 2,492.6
DCC-N 101.2 2,566.8
DCC-clik-QML 95.2 2,453.9
DCC-clik-N 97.7 2,528.9
DCC-t 101.5 2,555.3
DCC-AL 89.1 2,386.8
@@
To further backtest the VaR forecasts, we have employed several quantile accuracy and independence tests, including: the unconditional coverage (UC) test (Kupiec, 1995); the conditional coverage (CC) test (Christoffersen, 1998); the dynamic quantile (DQ) test (Engle and Manganelli, 2004); and the quantile regression based VaR (VQR) test (Gaglianone et al., 2011). Table 5 presents the UC, CC, DQ (lag 1 and lag 4) and VQR backtests’ p-values for the 10 competing models. The “Total” columns show the total number of rejections at the 5% significance level. As can be seen, the DCC-AL is the only framework which receives 0 rejections. The DCC-N model receives 5 rejections, while the DCC-t model gets rejected 3 times. This demonstrates the importance of the return distribution selection in parametric MGARCH models. The DCC-AL model, which is semi-parametric, shows advantage from this perspective. Comparing the DCC-AL model to the semi-parametric DCC models with QML, the DCC-AL still has better performance considering these backtests. These results again lend evidence on the validity and effectiveness of the DCC-AL in forecasting VaR.
@@
Table 5: Summary of 2.5% VaR forecasts UC, CC, DQ1, DQ4, and VQR backtests’ p-values and the total number of rejections at the 5% significance level.

### v4

@@ 4.1 Semi-DCC-AL performance
4.2 Standard DCC and Semi-DCC risk forecasting comparison
As mentioned, generating accurate portfolio risk forecasts via a semi-parametric approach is a key motivation when proposing the Semi-DCC framework. Therefore, to further demonstrate the effectiveness of the Semi-DCC in risk forecasting with the AL loss, the performance of Semi-DCC-AL is compared to the standard parametric DCC models, on the $1\%$, $2.5\%$ and $5\%$ probability levels. We first use the simulated datasets following the standardized multivariate Student’s $t$ distribution as in the previous section $\mathbf{z}_{t}\overset{iid}{\sim}t_{n}(\mathbf{0},I_{n};\nu)$ with the degrees of freedom parameter $\nu=10$. Then a heavier tailed choice of $\nu=5$ is also included. The Student’s $t$ is selected, as it can control the tail of the returns and the simulation results shown in the previous section with Normal and $nst$ are in general consistent with the ones from Student’s $t$.
The VaR and ES forecasting performance of the proposed Semi-DCC-AL framework on the 250 simulated datasets are compared to those yielded by the conventional DCC models fitted through 18 different DGPs which include: three probability levels $1\%$, $2.5\%$ and $5\%$ ; three different sample size $T\in\{2000,3000,5000\}$; two degrees of freedom parameters $\nu\in\{5,10\}$. For comparison, first we consider a standard parametric DCC-N and DCC-t models fitted by a two-stage procedure (Engle, 2002) combining maximization of Normal and Student’s $t$ distributions likelihoods in the first (volatility) and second (correlation) stages of the estimation procedure with the application of correlation targeting. For the VaR and ES forecasting, the theoretical quantile and tail expectation based on the Normal or Student’s $t$ distributions are used as detailed in the previous section. Further, the Composite Likelihood (clik) approach developed in Pakel et al. (2021) using Normal distribution is also included for comparison (DCC-clik-N).
[…]
    Across different probability levels, it can be seen that the more extreme 1% level creates more challenges for the parametric models, e.g., with $T=3000$, $\nu=5$, $1\%$ the ES forecasting RMSE values from parametric DCC models are more than doubled of the one from Semi-DCC-AL. On the least extreme 5% level, the performance gap of between parametric DCC and Semi-DCC-AL is smaller. Such findings confirm that the proposed semi-parametric approach is more robust on the more extreme probability level.
    Across the two different degrees of freedoms, the heavier tail choice of $\nu=5$ poses more challenges for all models, especially for the parametric DCC models. With $T=3000$ and $5\%$, when $\nu=10$ goes to $\nu=5$ the RMSE of ES forecast from parametric DCC models rises from $\approx 0.06$ to $\approx 0.11$ for parametric DCC models, while for Semi-DCC-AL the figure goes from 0.0579 to 0.0755. This demonstrates the usefulness of the semi-parametric approach in risk forecasting on the heavy tailed distributions, which are commonly observed in real financial time series data.
    Across different sample sizes, consistent with the study in the previous section, the RMSE of Semi-DCC-AL risk forecasts decreases as the time series length $T$ increases.
@@ 4.2 Standard DCC and Semi-DCC risk forecasting comparison
To conclude, the results of the simulation study support the two-step AL based estimation method which is able to produce accurate parameter estimates and portfolio risk forecasts, for both spherical and non-spherical error distributions. Comparing the risk forecasting results of the Semi-DCC-AL to several benchmarking parametric DCC models, the advantage of the proposed semi-parametric approach is clearly demonstrated, especially for the heavier-tailed return distribution and more extreme probably level.
Table 2: VaR forecasting results of parametric DCC models and Semi-DCC-AL with 250 simulation datasets. Red highlights RMSE values that are higher than the one from Semi-DCC-AL for each DGP.
DGP True-VaR DCC-N DCC-t DCC-clik-N Semi-DCC-AL
$t$, $T=2000$, $\nu=5$, $1\%$ Mean -1.7663 -1.5961 -1.6502 -1.5953 -1.7978
RMSE 0.1803 0.1537 0.1802 0.1802
$t$, $T=2000$, $\nu=5$, $2.5\%$ Mean -1.3545 -1.3450 -1.3669 -1.3443 -1.3792
RMSE 0.0285 0.0616 0.0282 0.1067
$t$, $T=2000$, $\nu=5$, $5\%$ Mean -1.0609 -1.1268 -1.1315 -1.1262 -1.0797
RMSE 0.0693 0.0812 0.0692 0.0691
$t$, $T=2000$, $\nu=10$, $1\%$ Mean -1.7766 -1.6839 -1.7215 -1.684 -1.7984
RMSE 0.1023 0.0998 0.1016 0.1374
$t$, $T=2000$, $\nu=10$, $2.5\%$ Mean -1.4361 -1.4180 -1.4253 -1.4181 -1.4499
RMSE 0.0292 0.0611 0.0287 0.0818
$t$, $T=2000$, $\nu=10$, $5\%$ Mean -1.1699 -1.1904 -1.1818 -1.1905 -1.1775
RMSE 0.0252 0.0464 0.0255 0.0552
$t$, $T=3000$, $\nu=5$, $1\%$ Mean -1.7729 -1.6007 -1.6522 -1.6003 -1.8002
RMSE 0.1784 0.1474 0.1786 0.1342
$t$, $T=3000$, $\nu=5$, $2.5\%$ Mean -1.3615 -1.3486 -1.3682 -1.3482 -1.3832
RMSE 0.0272 0.0537 0.0279 0.0775
$t$, $T=3000$, $\nu=5$, $5\%$ Mean -1.0682 -1.1318 -1.1341 -1.1315 -1.0881
RMSE 0.0669 0.0757 0.0671 0.0548
$t$, $T=3000$, $\nu=10$, $1\%$ Mean -1.7108 -1.6183 -1.6686 -1.6176 -1.7322
RMSE 0.0970 0.0834 0.0974 0.1056
$t$, $T=3000$, $\nu=10$, $2.5\%$ Mean -1.3825 -1.3634 -1.3817 -1.3629 -1.3990
RMSE 0.0252 0.0534 0.0256 0.0692
$t$, $T=3000$, $\nu=10$, $5\%$ Mean -1.1257 -1.1442 -1.1453 -1.1437 -1.1368
RMSE 0.0218 0.0454 0.0219 0.0469
$t$, $T=5000$, $\nu=5$, $1\%$ Mean -1.7736 -1.5985 -1.6544 -1.5977 -1.8067
RMSE 0.1857 0.1659 0.1860 0.1261
$t$, $T=5000$, $\nu=5$, $2.5\%$ Mean -1.3583 -1.3467 -1.3699 -1.3461 -1.3779
RMSE 0.0266 0.0649 0.0269 0.0724
$t$, $T=5000$, $\nu=5$, $5\%$ Mean -1.0658 -1.1302 -1.1355 -1.1297 -1.0825
RMSE 0.0697 0.0808 0.0695 0.0452
$t$, $T=5000$, $\nu=10$, $1\%$ Mean -1.7535 -1.6564 -1.6977 -1.6560 -1.7760
RMSE 0.1008 0.0907 0.1011 0.0892
$t$, $T=5000$, $\nu=10$, $2.5\%$ Mean -1.4155 -1.3955 -1.4058 -1.3952 -1.4304
RMSE 0.0232 0.0525 0.0235 0.0534
$t$, $T=5000$, $\nu=10$, $5\%$ Mean -1.1515 -1.1712 -1.1652 -1.1709 -1.1652
RMSE 0.0214 0.0415 0.0215 0.0387
Table 3: ES forecasting results of parametric DCC models and Semi-DCC-AL with 250 simulation datasets. Red highlights RMSE values that are higher than the one from Semi-DCC-AL for each DGP.
DGP True-ES DCC-N DCC-t DCC-clik-N Semi-DCC-AL
$t$, $T=2000$, $\nu=5$, $1\%$ Mean -2.3372 -1.8286 -1.9336 -1.8277 -2.3097
RMSE 0.5214 0.4337 0.5213 0.2613
$t$, $T=2000$, $\nu=5$, $2.5\%$ Mean -1.8555 -1.6043 -1.6672 -1.6035 -1.8663
RMSE 0.2578 0.2138 0.2578 0.1457
$t$, $T=2000$, $\nu=5$, $5\%$ Mean -1.5217 -1.4131 -1.4506 -1.4123 -1.5368
RMSE 0.1136 0.1005 0.1138 0.0963
$t$, $T=2000$, $\nu=10$, $1\%$ Mean -2.1620 -1.9291 -2.0173 -1.9293 -2.1411
RMSE 0.2412 0.1812 0.2404 0.1725
$t$, $T=2000$, $\nu=10$, $2.5\%$ Mean -1.8169 -1.6913 -1.7384 -1.6915 -1.8135
RMSE 0.1310 0.1145 0.1303 0.1045
$t$, $T=2000$, $\nu=10$, $5\%$ Mean -1.5546 -1.4929 -1.5153 -1.4930 -1.5558
RMSE 0.0659 0.0766 0.0653 0.0733
$t$, $T=3000$, $\nu=5$, $1\%$ Mean -2.3459 -1.8339 -1.9362 -1.8334 -2.3275
RMSE 0.5203 0.4310 0.5203 0.2024
$t$, $T=3000$, $\nu=5$, $2.5\%$ Mean -1.8652 -1.6086 -1.6689 -1.6081 -1.8778
RMSE 0.2616 0.2161 0.2617 0.1146
$t$, $T=3000$, $\nu=5$, $5\%$ Mean -1.5320 -1.4193 -1.4542 -1.4189 -1.5493
RMSE 0.1164 0.1016 0.1165 0.0755
$t$, $T=3000$, $\nu=10$, $1\%$ Mean -2.0819 -1.8540 -1.9555 -1.8532 -2.0688
RMSE 0.2328 0.1583 0.2332 0.1265
$t$, $T=3000$, $\nu=10$, $2.5\%$ Mean -1.7491 -1.6262 -1.6855 -1.6256 -1.7543
RMSE 0.1262 0.0977 0.1265 0.0850
$t$, $T=3000$, $\nu=10$, $5\%$ Mean -1.4958 -1.4349 -1.4686 -1.4343 -1.5021
RMSE 0.0637 0.0658 0.0641 0.0579
$t$, $T=5000$, $\nu=5$, $1\%$ Mean -2.3468 -1.8313 -1.9390 -1.8305 -2.3681
RMSE 0.5330 0.4513 0.5334 0.1949
$t$, $T=5000$, $\nu=5$, $2.5\%$ Mean -1.8608 -1.6063 -1.6712 -1.6056 -1.8824
RMSE 0.2616 0.2235 0.2620 0.1105
$t$, $T=5000$, $\nu=5$, $5\%$ Mean -1.5286 -1.4173 -1.4561 -1.4167 -1.5509
RMSE 0.1145 0.1084 0.1149 0.0671
$t$, $T=5000$, $\nu=10$, $1\%$ Mean -2.1339 -1.8977 -1.9899 -1.8972 -2.1370
RMSE 0.2406 0.1723 0.2409 0.1202
$t$, $T=5000$, $\nu=10$, $2.5\%$ Mean -1.7909 -1.6645 -1.7150 -1.6642 -1.7995
RMSE 0.1286 0.1043 0.1289 0.0703
$t$, $T=5000$, $\nu=10$, $5\%$ Mean -1.5301 -1.4687 -1.4942 -1.4684 -1.5415
RMSE 0.0628 0.0675 0.0631 0.0496
@@ 5 Empirical study
The performance of the Semi-DCC model in portfolio risk forecasting has been assessed via an empirical study on a multivariate time series of US stock returns. After providing a short description of the data and describing the forecasting design in Section 5.1, we assess the risk forecasting performance of the proposed models for an equally weighted portfolio in Section 5.2. This choice does not imply any loss of generality since our investigation could be easily replicated under alternative portfolio configurations. Our preference for the equally weighted scheme is motivated by the robust performance of this simple allocation rule that, in many instances, has been found to be competitive with more sophisticated benchmarks (DeMiguel et al., 2007). In addition, in Appendix A we also evaluate the effectiveness of the proposed framework in generating minimum risk portfolios, a task that cannot be accomplished with univariate risk forecasting models.
@@ 5.1 Equally weighted portfolio risk forecasting
Several existing univariate semi-parametric models are selected as benchmarks for comparison. First, we consider the recently proposed ES-CAViaR model (Taylor, 2019) with Symmetric Absolute Value (SAV) and Indirect GARCH (IG) specifications for the quantile regression equation. Regarding the ES specification, to facilitate the performance comparison with the Semi-DCC, we choose the model with the constant multiplicative ES to VaR factor $(1+\text{exp}(\gamma_{0}))$, which is the factor that we use in the proposed Semi-DCC model. Two VaR and ES joint loss functions, AL (Semi-DCC-AL) and FZ0 (Semi-DCC-FZ0), are employed and tested.
@@
For all the DCC benchmarks, in order to guarantee a fair comparison with the Semi-DCC model, the fitted univariate volatility specifications are given by GARCH(1,1) models.
In Figure (2), it is interesting to note that the estimated correlation parameters in the Semi-DCC-AL model vary over the forecasting period, reacting to changes in the underlying market volatility level. In particular, the estimated value of $a$ is characterized by a positive trend originating at the outbreak of the pandemic COVID-19 crisis. An opposite behaviour is observed for $b$.
Figure 2: Plots of estimated correlation parameters $a$ and $b$ of the Semi-DCC-AL model on the 2.5% probability level. The vertical lines denote the outbreak of the COVID-19 emergency (as officially declared by the World Health Organization on 30 January 2020). Whole forecasting sample (1213 days).
It is worth noting that the negative correlation emerges as an empirical property of the conditional correlation process. This behaviour is consistent with the recent literature, which provides evidence that conditional correlations have a state-dependent nature, driven mainly by the underlying volatility dynamics (see Bauwens and Otranto, 2016, among others). The link between the level of volatility and correlation dynamics is particularly evident in crisis periods, such as the one following the outbreak of the COVID 19 pandemics.
Table 4 compares the Semi-DCC-AL estimates (2.5% probability level) of correlation coefficients with those obtained for the benchmark DCC specifications considered, reporting the average estimates of correlation parameters $a$ and $b$ across all forecasting steps. Here, consistent with previous findings in the literature, standard DCC estimators tend to oversmooth (estimated $b$ very close to 1) correlations, while this is not the case for Semi-DCC-AL models. The DCC-clik estimates stay in between. It is here worth noting that the DCC-clik and Semi-DCC-AL models are based on different first stage volatility estimators: QML-GARCH for DCC-clik, and ES-CAViaR-IG as in Equation (7) for Semi-DCC-AL. In addition, Figure 3 shows the cross-sectional averages of the estimated conditional correlations (calculated based on $P_{t}$ in Equation (12)) across the forecasting period. As can be seen, the dynamics of the estimated conditional correlations between Semi-DCC-AL and DCC-clik are relatively close to each other.
Table 4: Average estimates (across time) of correlation dynamic parameters ($a$ and $b$) from different models and estimation methods on the 2.5% probability level. Whole forecasting sample (1213 days).
Semi-DCC-AL DCC-QML DCC-t DCC-clik
@@
Figure 3: Cross-sectional averages of the estimated conditional correlation for DCC-QML, DCC-t, DCC-clik and Semi-DCC-AL on the 2.5% probability level, across the whole forecasting sample (1213 days).
The average of estimated first stage univariate modeling coefficients on the 2.5% probability level, across all assets and forecasting steps, is reported in Table 5. GARCH-$t$ and QML-GARCH, used in DCC-t and DCC-QML respectively, appear to be less reactive (smaller $\alpha$ estimates) to past shocks compared to the ES-CAViaR-IG, used in the Semi-DCC-AL model.
Table 5: Average estimates (across times and assets) of the first stage univariate modeling parameters ($\alpha$ and $\beta$) from different models and estimation methods on the 2.5% probability level, across the full forecasting sample (1213 days). For ES-CAVIAR-IG, we report $\alpha=\alpha^{(q)}/q^{2}$.
@@ 5.2 Evaluation of VaR and ES forecasts
Assuming equally weighted portfolio returns, one-step-ahead daily VaR and ES forecasts from the proposed Semi-DCC-AL and Semi-DCC-FZ0 models are produced by using Equation (15). In this section, these forecasts are compared with the ones from the competing models presented in Section 5.1.
@@
$L_{Q}=\sum_{t=T_{\text{in}}+1}^{T_{\text{in}}+T_{\text{out}}}(r_{t}-\widehat{Q}_{t})(\alpha-I(r_{t}\leq\widehat{Q}_{t}))\,\,,$ (19)
@@
$L_{\text{AL},t}(r_{t},\widehat{Q}_{t},\widehat{\text{ES}}_{t})=-\text{log}\left(\frac{\alpha-1}{\widehat{\text{ES}}_{t}}\right)-{\frac{(r_{t}-\widehat{Q}_{t})(\alpha-I(r_{t}\leq\widehat{Q}_{t}))}{\alpha\widehat{\text{ES}}_{t}}}.$ (20)
@@
In our analysis, we first use the joint loss $L_{\text{AL}}=\sum_{t=T_{\text{in}}+1}^{T_{\text{in}}+T_{\text{out}}}L_{\text{AL},t}$ to formally and jointly assess and compare the VaR and ES forecasts from all models.
Meanwhile, the FZ0 loss used in Patton et al. (2019) can be calculated as:
$L_{\text{FZ0},t}(r_{t},\widehat{Q}_{t},\widehat{\text{ES}}_{t})=-\frac{1}{\alpha\text{ES}_{t}}I(r_{t}\leq Q_{t})\,(Q_{t}-r_{t})+\frac{Q_{t}}{\text{ES}_{t}}+\log(-\text{ES}_{t})-1.$ (21)
The FZ0 joint loss $L_{\text{FZ0}}=\sum_{t=T_{\text{in}}+1}^{T_{\text{in}}+T_{\text{out}}}L_{\text{FZ0},t}$ is also used to test the performance of the models in comparison.
@@
Figure 4 visualizes the 2.5% portfolio ES forecasts from the DCC-QML, DCC-clik-QML and Semi-DCC-AL models for the Jan 2019–Sep 2021 period, using equally weighted Dow Jones returns. In general, we can see that the ES forecasts produced from the Semi-DCC-AL model have a long-run behaviour comparable to the ones from the DCC-QML and DCC-clik-QML. However, in the short term, forecasts from these models can be characterized by substantially different dynamic patterns. This is particularly evident in the period immediately following the outbreak of the COVID-19.
Inspecting the ES forecasts at the beginning of 2020, we can see that the three models have distinctive behaviours. Comparing to the DCC-QML model, the Semi-DCC-AL model is more reactive to the return shocks. This is because of the larger $a$ and $\alpha$ estimates as discussed in Tables 4 and 5. The ES forecasts from the DCC-clik-QML stay in between the ones from DCC-QML and Semi-DCC-AL, which is also consistent with the findings from Table 4.
@@
Figure 4: 2.5% ES forecasts from the DCC-QML, DCC-clik-QML and Semi-DCC-AL models, using equally weighted Dow Jones returns.
@@
The quantile and joint loss results from the proposed Semi-DCC-AL and Semi-DC-FZ0 models and other competing models are shown in Table 6. Overall, the results show that the proposed Semi-DCC-AL and Semi-DCC-FZ0 models, in comparison to the other models under analysis, generate competitive loss results, which lends evidence on the validity of the proposed semi-parametric framework. Furthermore, the Semi-DCC models generate smaller loss values than all other parametric or semi-parametric DCC models that have been taken as benchmarks. Meanwhile, the performance of Semi-DCC-AL and Semi-DCC-FZ0 is quite close to each other, showing that the choice of the joint loss has a minimal impact on the Semi-DCC estimation and risk forecasting performance.
@@
The Model Confidence Set (MCS) of Hansen et al. (2011) is employed to statistically compare the quantile loss (Equation (19)), AL joint loss (Equation (20)) and FZ0 joint loss (Equation (21)) values yielded by the different models. A MCS is a set of models constructed to contain the superior models with a given level of confidence, chosen in our paper as 75%. The SQ method (see Hansen et al., 2011, for details) is employed to calculate the MCS test statistic. The MCS results, using both quantile and AL&FZ0 joint loss functions, are also included in Table 6, with red highlighting indicates models that are excluded from the 75% MCS. First, the AL and FZ0 loss function produce consistent MCS results as expected. Further, as can be seen the Semi-DCC-AL and Semi-DCC-FZ0 are always included in the MCS for the quantile test and both joint loss tests, across all probability levels. On the more extreme 1% level, for the quantile loss test only CARE-IG is in the MCS in addition to the Semi-DCC. For the joint loss tests, most of the existing DCC models are not in the MCS, except DCC-clik-QML. On the 2.5% level, all the DCC models, except the proposed Semi-DCC, are excluded from the MCS based on the joint loss tests. On the 5% level, only the DCC-clik-N and Semi-DCC models are in the MCS based on joint loss.
@@
Table 6: The quantile loss (Equation (19)), AL joint loss (Equation (20)) and FZ0 joint loss values (Equation (21)) for all competing models across the full forecasting sample. Models in red are excluded from the 75% MCS according to the SQ statistic.
Model Quantile loss AL Joint loss FZ0 joint loss
$\bm{\alpha=1\%}$
CARE-SAV 43.8 2,741.2 1,491.4
CARE-IG 39.3 2,540.4 1,294.9
ES-CAViaR-SAV 40.5 2,576.7 1,331.4
ES-CAViaR-IG 40.0 2,547.8 1,301.7
DCC-QML 41.8 2,628.8 1,385.6
DCC-N 43.8 2,720.3 1,472.1
DCC-clik-QML 40.6 2,585.5 1,340.7
DCC-clik-N 42.1 2,679.5 1,429.6
DCC-t 43.6 2,695.1 1,447.4
Semi-DCC-AL 38.9 2,528.5 1,282.6
Semi-DCC-FZ0 38.9 2,528.3 1,282.5
$\bm{\alpha=2.5\%}$
CARE-SAV 90.4 2,405.4 1,133.8
CARE-IG 89.1 2,392.1 1,120.4
ES-CAViaR-SAV 89.4 2,398.1 1,126.5
ES-CAViaR-IG 87.3 2,369.5 1,097.4
DCC-QML 98.7 2,492.6 1,221.8
DCC-N 101.2 2,566.8 1,290.8
DCC-clik-QML 95.2 2,453.9 1,181.4
DCC-clik-N 97.7 2,528.9 1,251.3
DCC-t 101.5 2,555.3 1,279.8
Semi-DCC-AL 89.1 2,386.8 1,115.0
Semi-DCC-FZ0 89.0 2,383.8 1112.0
$\bm{\alpha=5\%}$
CARE-SAV 130.0 2,070.2 764.9
CARE-IG 130.0 2,070.7 765.9
ES-CAViaR-SAV 131.3 2,077.4 770.1
ES-CAViaR-IG 130.6 2,065.1 757.9
DCC-QML 137.2 2,157.3 856.2
DCC-N 136.2 2,143.9 839.0
DCC-clik-QML 134.1 2,123.3 820.1
DCC-clik-N 132.7 2,108.4 801.2
DCC-t 136.5 2,138.5 833.4
Semi-DCC-AL 131.1 2,079.9 772.9
Semi-DCC-FZ0 131.1 2,083.9 776.7
@@
To further backtest the VaR forecasts, we have employed several quantile accuracy and independence tests, including: the unconditional coverage (UC) test (Kupiec, 1995); the conditional coverage (CC) test (Christoffersen, 1998); the dynamic quantile (DQ) test (Engle and Manganelli, 2004); and the quantile regression based VaR (VQR) test (Gaglianone et al., 2011). Table 7 presents the UC, CC, DQ (lag 1 and lag 4) and VQR backtests’ p-values for the 11 competing models. The “Total” columns show the total number of rejections at the 5% significance level. As can be seen, across the 1% and 2.5% probability levels the Semi-DCC is the only framework which receives 0 rejections. On the 5% level, the Semi-DCC is rejected 3 times, while all the other models are rejected by all the 5 tests. The DCC-N and DCC-t models are more likely to be rejected by various tests. This demonstrates the importance of the return distribution selection in parametric DCC models. The Semi-DCC model, which is semi-parametric, shows advantage from this perspective. Comparing the Semi-DCC model to the other semi-parametric DCC models with QML, the Semi-DCC still has better performance considering these backtests. These results again lend evidence on the validity and effectiveness of the Semi-DCC in VaR forecasting.
@@
Table 7: Summary of VaR forecasts UC, CC, DQ1, DQ4, and VQR backtests’ p-values and the total number of rejections at the 5% significance level.
@@
$\bm{\alpha=1\%}$
CARE-SAV 2% 6% 23% 63% 1% 2
CARE-IG 34% 60% 78% 98% 64% 0
ES-CAViaR-SAV 20% 42% 69% 4% 76% 1
ES-CAViaR-IG 34% 60% 78% 98% 60% 0
DCC-QML 53% 15% 3% 0% 82% 2
DCC-N 19% 3% 0% 0% 6% 3
DCC-clik-QML 97% 2% 0% 0% 89% 3
DCC-clik-N 7% 2% 0% 0% 0% 4
DCC-t 19% 3% 0% 0% 5% 3
Semi-DCC-AL 20% 42% 67% 95% 78% 0
Semi-DCC-FZ0 20% 42% 67% 95% 65% 0
$\bm{\alpha=2.5\%}$